January 11, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, January 11, 2017
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates
  3. Data Validation
  4. GlueX Detector Performance Document

Upcoming Calibrations

  • Add timing information to FCAL clustering/inner ring calibrations? (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • TAGH alignment (Nathan)
  • Tagger energy map (Richard J.?)
  • TOF alignment (Beni/Simon)


  1. Announcements
  2. Calibration Updates
    • Calibrations train
      • Several iterations run, including TAGH and TOF improvements
      • Results from full runs available:
        • Skims: /cache/halld/RunPeriod-2016-10/calib
        • Histograms: /cache/halld/RunPeriod-2016-10/calib/hists/summed
      • Timing alignment is generally okay, drift chambers needed attentions as higher noise on CDC and FDC sometimes threw off fits, effect of new TOF calibrations from Beni last Thursday need to be studied.
      • Waiting on monitoring run to assess current situation. Current recon. test results.
    • Subdetectors
    • Alignment
  3. Simulations
    • sim1.2 progress
  4. AOB


Attendees: Alex A., Thomas B., Brad C., Eugene C., Mark D., Sergey F., Nacer H., Mark I., Mahmoud K., Paul M., Curtis M., Lubomir M., Matt S., Nathan S., Mike S., Justin S., George V.,

  1. Thomas B. showed us the brand-new calibration web-browser interface. All the fields except the "step number" are auto-generated.
  2. Brad has been working on analysing the TOF HV scan data but his efforts have been hampered by Lustre problems.
    1. Computer center is working on the Lustre issues...
  3. Data from the Fall'16 run have been processed on the farm and REST files have been produced. Initial calibrations were performed by Sean, but refinement by the various detector groups may be needed. Some early runs may not be calibratable because of TDC initialization problems.
  4. Mike S. is continuing to work on the FDC/CDC aligment. He has had some success using millepede to determine parameters for the FDC cathodes (resolution went from 200 microns to 155 microns) but the FDC anodes have proven to be trickier. The code for the CDC straight-line fitter needed work -- the smoother was not implemented yet. All the work on the code is on a branch on github -- merging into the main trunk will take place at a later date.
  5. Mark I. started sim1.2 over the break; ~6300 jobs have run and many more are currently pending on the farm. The code base is not quite consistent with REST recon3. It also appears to be running more slowly than previous simulation passes, with more time-out failures. Not sure of origin of this.